given the equation of an ellipse,9x^2+4y^2+36x-8y+4=0,Please show the solution and answer
Answer by richwmiller(17219)

9x^2+4y^2+36x-8y+4=0
there is no solution because there are two variables nor answer because there is no question.
ellipse
foci | ((-2, 1-sqrt(5)) | (-2, 1+sqrt(5)))=((-2, -1.23607) | (-2, 3.23607))
vertices | (-2, -2) | (-2, 4)
center | (-2, 1)
semimajor axis length | 3
semiminor axis length | 2
area | 6 pi~~18.8496
perimeter | 12 E(5/9)=15.8654
focal parameter | 4/sqrt(5)=1.78885
eccentricity | sqrt(5)/3=0.745356